# Molecular-type $QQss\bar{s}$ pentaquarks predicted by the extended chiral unitary approach

### 26 Dec 2023 v2

@@ III Numerical results
Next, in the $bbss\bar{s}$ sector, we show our results in Fig. 3. From the modulus square of the amplitudes we can see four distinct peaks around the energy range $11380$ MeV to $11480$ MeV. The poles and their coupling constants for each coupled channel are given in Table 8. The four states obtained are the $\bar{B}_{s}\Omega_{b}$, $\bar{B}_{s}\Omega_{b}^{*}$, $\bar{B}_{s}^{*}\Omega_{b}$, and $\bar{B}_{s}^{*}\Omega_{b}^{*}$ bound states, respectively, which are all found on the first Riemann sheet. Due to the absence of the exchange of light pseudoscalar and $B_{s}^{*}$ mesons, which means that there are no coupled channels for each other, see the coefficients in Table 4, the obtained poles are all without imaginary parts. The thresholds of the $\bar{B}_{s}\Omega_{b}$, $\bar{B}_{s}\Omega_{b}^{*}$, $\bar{B}_{s}^{*}\Omega_{b}$, and $\bar{B}_{s}^{*}\Omega_{b}^{*}$ channels are $11412.12$ MeV, $11435.92$ MeV, $11460.60$ MeV, and $11484.40$ MeV, respectively, see Table 6, and so we have these states all bound at about $21$ MeV.
@@
Note that for the molecular states obtained above, the attractions of their most relevant channels are provided by the exchange of the vector meson $\phi$, while the interactions in nondiagonal terms are weaker due to the exchange of the heavy vector meson $D_{s}^{*}$. It is very interesting that in the absence of long-range interactions mediated by the exchange of $\pi$ meson, the bound states can only be formed by the exchange of the heavier meson $\phi$. These types of doubly heavy pentaquarks with the most strange quark contents $ccss\bar{s}$, $bbss\bar{s}$, and $bcss\bar{s}$ are excellent systems to validate this mechanism. We look forward to the experimental search for these states to confirm the model used in this paper. If these states exist, it will deepen our understanding of the nonperturbative behavior of the strong interaction.
@@ IV Summary
In the $ccss\bar{s}$ sector, we obtain four bound states mainly coupled to the $D_{s}\Omega_{c}$, $D_{s}\Omega_{c}^{*}$, $D_{s}^{*}\Omega_{c}$, and $D_{s}^{*}\Omega_{c}^{*}$ channels, respectively. In the $bbss\bar{s}$ sector, we obtain four bound states mainly coupled to the $\bar{B}_{s}\Omega_{b}$, $\bar{B}_{s}\Omega_{b}^{*}$, $\bar{B}_{s}^{*}\Omega_{b}$, and $\bar{B}_{s}^{*}\Omega_{b}^{*}$ channels, respectively. In the $bcss\bar{s}$ sector, we obtain eight bound states mainly coupled to $D_{s}\Omega_{b}$, $\bar{B}_{s}\Omega_{c}$, $D_{s}\Omega_{b}^{*}$, $\bar{B}_{s}\Omega_{c}^{*}$, $D_{s}^{*}\Omega_{b}$, $\bar{B}_{s}^{*}\Omega_{c}$, $D_{s}^{*}\Omega_{b}^{*}$, and $\bar{B}_{s}^{*}\Omega_{c}^{*}$ channels, respectively. In the cases of the $PB(1/2^{+})$, $PB(3/2^{+})$, $VB(1/2^{+})$, and $VB(3/2^{+})$ sectors, which carry the quantum numbers $I(J^{P})=0(1/2^{-})$, $0(3/2^{-})$, $0(1/2^{-},3/2^{-})$, and $0(1/2^{-},3/2^{-},5/2^{-})$, respectively. All these states have very small or zero widths, which are below the threshold of the most dominant channel about $20-30$ MeV. The LHCb Collaboration is carrying out the relevant measurement and analysis, and these predicted states are likely to be detected in the near future.
@@ Acknowledgements
We would like to thank Fu-Lai Wang and Wen-Fei Wang for valuable discussions, and acknowledge Profs. Eulogio Oset and Kazem Azizi for careful reading the manuscript and useful comments. This work is supported by the China National Funds for Distinguished Young Scientists under Grant No. 11825503, the National Key Research and Development Program of China under Contract No. 2020YFA0406400, the 111 Project under Grant No. B20063, the fundamental Research Funds for the Central Universities under Grant No. lzujbky-2022-sp02, the project for top-notch innovative talents of Gansu province, and the National Natural Science Foundation of China (NSFC) under Grants No. 12247101, 12335001, 11965016, 11705069 and 12047501 (Z.F.S.). This work is also partly supported by the Natural Science Foundation of Changsha under Grants No. kq2208257, the Natural Science Foundation of Hunan province under Grant No. 2023JJ30647, and the NSFC under Grant No. 12365019 (C.W.X.).

### 08 Feb 2024 v4

# Molecular-type $QQss\bar{s}$ pentaquarks predicted by an extended hidden gauge symmetry approach

@@ III Numerical results
Next, in the $bbss\bar{s}$ sector, we show our results in Fig. 3. From the modulus square of the amplitudes we can see four distinct peaks around the energy range $11380$ MeV to $11480$ MeV. The poles and their coupling constants for each coupled channel are given in Table 8. The four states obtained are the $\bar{B}_{s}\Omega_{b}$, $\bar{B}_{s}\Omega_{b}^{*}$, $\bar{B}_{s}^{*}\Omega_{b}$, and $\bar{B}_{s}^{*}\Omega_{b}^{*}$ bound states, respectively, which are all found on the first Riemann sheet. Due to the absence of the exchange of light pseudoscalar and $B_{s}^{*}$ mesons, which means that there are no coupled channels for each other, see the coefficients in Table 4, the obtained poles are all without imaginary parts and all the couplings of the lower channels are zero. This is different from the case of the $ccss\bar{s}$ sector with the $D_{s}^{*}$ exchange, see the results in Table 7. The thresholds of the $\bar{B}_{s}\Omega_{b}$, $\bar{B}_{s}\Omega_{b}^{*}$, $\bar{B}_{s}^{*}\Omega_{b}$, and $\bar{B}_{s}^{*}\Omega_{b}^{*}$ channels are $11412.12$ MeV, $11435.92$ MeV, $11460.60$ MeV, and $11484.40$ MeV, respectively, see Table 6, and thus we have these states all bound at about $21$ MeV, because of in fact the similar single channel interaction.
@@
Note that for the molecular states obtained above, the attractions of their most relevant channels are provided by the exchange of the vector meson $\phi$, while the interactions in nondiagonal terms are weaker due to the exchange of the heavy vector meson $D_{s}^{*}$. It is very interesting that in the absence of long-range interactions mediated by the exchange of $\pi$ meson, the bound states can only be formed by the exchange of the heavier meson $\phi$. Thus, the binding energies of these systems are only tens MeV, which become more reasonable in the molecular picture for the heavy sector. It should be mentioned that in some cases, the deep bound systems are also found, such as the $\Lambda_{c}(2595)$ state in the $DN$ system Hofmann:2005sw ; Mizutani:2006vq and the $f_{2}(1270)$ in the $\rho\rho$ interaction Molina:2008jw . Therefore, these types of doubly heavy pentaquarks with the most strange quark contents $ccss\bar{s}$, $bbss\bar{s}$, and $bcss\bar{s}$ are excellent systems to validate this mechanism. We look forward to the experimental search for these states to confirm the model used in this paper. If these states exist, it will deepen our understanding of the nonperturbative behavior of the strong interaction.
@@ IV Summary
In the $ccss\bar{s}$ sector, we obtain four bound states mainly coupled to the $D_{s}\Omega_{c}$, $D_{s}\Omega_{c}^{*}$, $D_{s}^{*}\Omega_{c}$, and $D_{s}^{*}\Omega_{c}^{*}$ channels, respectively. In the $bbss\bar{s}$ sector, we obtain four bound states mainly coupled to the $\bar{B}_{s}\Omega_{b}$, $\bar{B}_{s}\Omega_{b}^{*}$, $\bar{B}_{s}^{*}\Omega_{b}$, and $\bar{B}_{s}^{*}\Omega_{b}^{*}$ channels, respectively. In the $bcss\bar{s}$ sector, we obtain eight bound states mainly coupled to $D_{s}\Omega_{b}$, $\bar{B}_{s}\Omega_{c}$, $D_{s}\Omega_{b}^{*}$, $\bar{B}_{s}\Omega_{c}^{*}$, $D_{s}^{*}\Omega_{b}$, $\bar{B}_{s}^{*}\Omega_{c}$, $D_{s}^{*}\Omega_{b}^{*}$, and $\bar{B}_{s}^{*}\Omega_{c}^{*}$ channels, respectively. In the cases of the $PB(1/2^{+})$, $PB(3/2^{+})$, $VB(1/2^{+})$, and $VB(3/2^{+})$ sectors, which carry the quantum numbers $I(J^{P})=0(1/2^{-})$, $0(3/2^{-})$, $0(1/2^{-},3/2^{-})$, and $0(1/2^{-},3/2^{-},5/2^{-})$, respectively. All these states have very small or zero widths, which are below the threshold of the most dominant channel about $20-30$ MeV. Some of these predicted molecular pentaquark candidates can be searched in the corresponding decay channels $\eta\Omega_{cc}^{(*)}$, $\phi\Omega_{cc}^{(*)}$, $\eta^{\prime}\Omega_{bc}^{\prime/*}$, $\phi\Omega_{bc}^{\prime/*}$, and so on. The LHCb Collaboration is carrying out the relevant measurement and analysis, and these predicted states are likely to be detected in the near future.
@@ Acknowledgements
We would like to thank Fu-Lai Wang and Wen-Fei Wang for valuable discussions, and acknowledge Profs. Eulogio Oset and Kazem Azizi for careful reading the manuscript and useful comments. This work is supported by the China National Funds for Distinguished Young Scientists under Grant No. 11825503, the National Key Research and Development Program of China under Contract No. 2020YFA0406400, the 111 Project under Grant No. B20063, the fundamental Research Funds for the Central Universities under Grant No. lzujbky-2022-sp02, the project for top-notch innovative talents of Gansu province, and the National Natural Science Foundation of China (NSFC) under Grants No. 12247101, 12335001, 11965016, 11705069 and 12047501 (Z.F.S.). This work is also partly supported by the Natural Science Foundation of Changsha under Grants No. kq2208257, the Natural Science Foundation of Hunan province under Grant No. 2023JJ30647, the Natural Science Foundation of Guangxi province under Grant No. 2023JJA110076, and the NSFC under Grant No. 12365019 (C.W.X.).